OpenStudy (anonymous):

**PLEASE HELP! FAN AND MEDAL!!!** 4/5x + 4/3 = 2x solve for X

OpenStudy (anonymous):

\[4/5x + 4/3 = 2x\]

OpenStudy (darkbluechocobo):

You do realize this is history not math XD

OpenStudy (darkbluechocobo):

but i shall help

OpenStudy (anonymous):

4/5x + 4/3 = 2x - 1 1/5x = - 4/3 6/5x = 4/3 (5/6)6/5x = (5/6)4/3 x = 10/9 or x = 1 1/9.

OpenStudy (anonymous):

sorry I didn't see that, usualy when I log on im in the math section XD sorry

OpenStudy (darkbluechocobo):

XD